# Towards Multi-Granularity Memory Association and Selection for Long-Term Conversational Agents

### v1

@@ 3.2 Overall Results
We present the results for Question Answering and Retrieval in Table 1 and Table 2, respectively. We also compare performence of single-granularity and multi-granularity in Table 3. Note that LongMTBench+ does not provide a ground-truth retrieval corpus and is therefore not included. RAPTOR cannot be evaluated for retrieval performance because its retrieved information cannot be judged. Due to high runtime of RAPROT and HippoRAG on LongMemEval-m, we do not obtain results on them. Below, we provide analysis of these results.
Table 1: QA Performance. Contriever is used as the retrieval backbone (except for Full History and MPNet), with GPT4o-mini as the generator, and the top 3 results are retrieved as input context. Bolded values indicate the best results, while underlined values represent the second-best results. All metrics are the larger, the better.
Model GPT4o-J F1 BLEU Rouge1 Rouge2 RougeL BERTScore
LongMemEval-s
Full History 50.60 11.48 1.40 12.10 5.47 10.85 83.07
MPNet [43] 53.20 13.96 2.21 14.49 6.78 12.93 83.72
Contriever [13] 55.40 13.78 2.21 14.46 6.93 12.89 83.70
LLM-Rsum [49] 35.40 12.29 2.09 13.01 5.55 11.52 83.60
MPC [19] 53.80 13.60 1.74 14.27 6.49 12.95 83.49
SeCom [33] 56.00 12.95 2.25 13.80 6.09 11.93 83.51
HippoRAG 2[9] 57.60 14.73 2.15 15.30 7.36 13.83 83.86
RAPTOR [42] 32.20 12.08 1.90 12.73 5.82 11.25 83.50
MemGAS (Ours) 60.20 20.38 4.22 21.05 10.47 19.47 85.21
LongMemEval-m
Full History 12.20 5.70 0.78 6.27 2.08 5.28 81.62
MPNet [43] 37.80 10.76 1.70 11.46 4.76 10.03 83.09
Contriever [13] 42.80 11.88 1.66 12.56 5.63 11.02 83.28
LLM-Rsum [49] 23.80 10.04 1.70 10.89 4.26 9.21 83.12
MPC [19] 37.80 11.28 1.37 11.93 5.12 10.57 82.98
SeCom [33] 42.80 11.33 1.79 12.03 5.07 10.49 83.16
MemGAS (Ours) 45.40 16.85 3.39 17.60 8.25 16.14 84.69
LOCOMO-10
Full History 33.43 12.23 1.84 12.70 5.66 11.73 84.07
MPNet [43] 38.07 14.44 2.35 14.90 6.83 13.90 84.42
Contriever [13] 40.33 15.66 2.67 16.01 7.68 15.00 84.65
LLM-Rsum [49] 22.56 9.14 0.99 9.82 3.38 8.98 83.45
MPC [19] 40.38 14.81 1.99 15.10 6.83 14.13 84.43
SeCom [33] 44.21 13.79 2.30 14.28 6.17 13.30 84.04
HippoRAG 2[9] 45.62 16.66 2.91 17.01 8.27 15.93 84.88
RAPTOR [42] 31.72 14.55 2.88 15.09 7.49 14.18 84.48
MemGAS (Ours) 41.07 17.66 3.61 18.00 8.93 16.99 85.13
LongMTBench+
Full History 67.44 36.07 11.32 37.90 20.51 29.25 87.81
MPNet [43] 63.89 36.09 11.26 38.39 20.58 28.86 87.84
Contriever [13] 63.54 36.30 11.59 38.17 21.65 29.67 87.90
LLM-Rsum [49] 24.65 26.58 6.91 29.23 11.93 20.90 86.11
MPC [19] 61.81 31.52 7.97 33.56 17.27 25.20 86.51
SeCom [33] 64.58 36.68 12.01 38.81 21.44 29.65 87.88
HippoRAG 2[9] 63.54 35.64 11.05 37.61 20.37 28.76 87.70
RAPTOR [42] 59.72 37.69 13.47 40.08 21.68 30.88 88.38
MemGAS (Ours) 69.44 41.49 15.62 43.69 24.45 34.66 88.96
Question Answering Results. As presented in Table 1, MemGAS consistently outperforms other methods across most datasets and evaluation metrics. Unlike the Full History approach, which introduces noise by utilizing all historical information, MemGAS excels by effectively consolidating and retrieving only the most relevant memories. Other baselines, such as LLM-Rsum and SeCom, although operating at specific granularities, lack the capability to integrate multi-granular associations between memory, resulting in suboptimal outcomes. Besides, methods like HippoRAG 2 and RAPTOR, while establishing connections between memory units, fail to construct multi-granular relationships and selection mechanisms effectively, limiting their performance. MemGAS performs better through its dynamic construction and adaptive router of multi-granular memory units and redundancy filtering, emphasizing the critical role of association and selection in effective long-term memory management.
@@
MPNet [43] 66.17 75.47 76.38 78.29 85.11 80.63
Contriever [13] 71.06 79.72 81.28 82.47 90.00 84.29
LLM-Rsum [49] 67.23 78.33 79.79 81.76 87.66 83.28
MPC [19] 60.00 70.90 68.09 73.27 80.00 76.59
SeCom [33] 71.06 80.88 80.43 83.08 89.15 85.11
HippoRAG 2[9] 75.53 85.44 84.68 87.32 91.28 88.73
@@
MPNet [43] 37.02 48.68 45.74 52.51 61.28 56.55
Contriever [13] 44.26 56.11 53.40 59.44 65.96 62.74
LLM-Rsum [49] 23.19 32.10 31.70 36.99 43.62 41.40
MPC [19] 35.96 47.51 42.55 50.36 54.26 53.88
SeCom [33] 44.26 56.61 55.32 60.76 66.60 63.78
@@
LOCOMO-10
MPNet [43] 45.92 47.71 53.98 51.79 68.58 56.88
Contriever [13] 49.90 52.15 58.26 56.29 71.80 60.92
LLM-Rsum [49] 47.23 48.99 59.01 54.58 74.97 60.07
MPC [19] 49.50 51.47 57.45 55.53 71.85 60.47
SeCom [33] 55.24 57.90 64.80 62.36 78.30 66.97
HippoRAG 2[9] 56.60 58.37 65.06 62.50 78.05 66.79
@@
Table 3: Single-Granularity vs Multi-Granularity on QA performance, which is evaluated using top 3 retrieved results.
Granularity Method QA Performance Retrieval Performance
GPT4o-J F1 RogueL Recall@3 Recall@5 Recall@10
Single- Granularity Session-level 55.40 13.78 12.89 71.06 81.28 90.00
Turn-level 57.60 14.94 14.12 73.62 84.68 91.91
Summary-level 28.80 10.66 9.96 70.43 80.00 88.30
Keyword-level 17.20 9.05 8.20 62.98 74.68 82.34
Multi- Granularity Combination 56.60 14.59 13.70 76.17 87.23 91.91
MemGAS 60.20 20.38 19.47 78.51 88.94 94.47
Single-granularity vs Multi-granularity Results. We compare our MemGAS with Single-granularity settings including (session-level, turn-level, summary-level and keyword-level), and Multi-granularity by directly combining text. Table 3 shows that single-granularity settings underperform compared to multi-granularity approaches in both QA and retrieval tasks. This demonstrates the effectiveness of integrating multi-granular information in the long-term memory of agents. Furthermore, MemGAS achieves the best results, surpassing the simple combination of multi-granularity text. This indicates that the proposed memory association and entropy-based router modules effectively utilize multi-granular information, leading to significant performance gains. Comprehensive analyses across datasets and metrics further support these findings (See in Appendix B).
@@ 3.3 Ablation Study
The ablation study in Table 4 illustrates the contributions of various components to the QA and retrieval performance of MemGAS. In the w/o Gaussian Mixture Model setting, the system links using only the most similar node, resulting in a noticeable performance drop. Similarly, the w/o Personalized PageRank setting, which uses one-hot nodes retrieval, leads to reduced recall and QA scores, emphasizing the importance of multi-hop retrieval. Memory Association and Router also contribute significantly, with their absence leading to declines in QA and retrieval results. Interestingly, removing filtering boosts GPT4o-J’s QA performance. It can be explained that retaining more information aligns better with the LLM judge’s preference for comprehensive responses. The most significant drop occurs when all components are removed, with the F1 score decreasing from 20.38 to 13.78, highlighting their combined importance.
Table 4: Ablation Study on the LongMemeval-s Dataset. QA performance is evaluated using top 3 retrieved results. The filtering setting is applied as a post-retrieval process. Thus, retrieval results are not included.
@@
GPT4o-J F1 RogueL Recall@3 Recall@5 Recall@10
MemGAS 60.20 20.38 19.47 78.51 88.94 94.47
w/o Gaussian Mixture Model 57.20 19.49 18.68 76.38 85.53 91.28
w/o Personalized PageRank 56.60 19.76 18.85 75.96 85.96 90.64
w/o Memory Association 56.80 17.69 19.00 74.89 85.74 91.49
w/o Soft Router 56.60 18.88 18.62 75.53 85.74 92.34
w/o Filtering 60.60 14.37 13.54 - - -
w/o All 55.40 13.78 12.89 71.06 81.28 90.00
@@
3.4 Detailed Comparison Analysis
Comparison on different query types. The results in Figure 6 highlight the performance across different query types. Our MemGAS consistently demonstrates superior performance, particularly excelling in multi-hop retrieval on Locomo-10 and multi-session on LongMemEval-s. This suggests that the memory association mechanism in MemGAS effectively identifies highly relevant sessions, enabling enhanced multi-hop reasoning and better integration of knowledge. Additionally, our method performs well across single-session, temporal reasoning, and knowledge update, demonstrating its strength in addressing both simple and complex query types. We also provide a case study for different query types in Appendix I.1, and comparison with other datasets and metrics in Appendix E.
Comparison on different Top-K retrieval. The results shown in Figure 4 demonstrate the performance of various Top-K retrieval settings. On the LoCoMo-10 and LongMemEval-s datasets, our model consistently surpasses baseline methods in F1 scores as Top-K increases from 1 to 10. This trend highlights that retrieving more relevant context significantly enhances accuracy. Note that on LongMemEval-s dataset, while F1 scores initially improve at lower Top-K values, performance declines at higher Top-K levels, suggesting that the longer context introduces noise, which negatively impacts the model’s effectiveness.
@@ 4 Related Work
Retrieval Methods. Recent advancements in retrieval methods have greatly enhanced information retrieval performance[40, 43, 41]. BM25 [40], a classic sparse retrieval method, uses a probabilistic model to improve query relevance. In contrast, dense retrieval methods excel at capturing semantic similarity. For instance, DPR [16] employs dual-encoder models to encode queries and documents into dense vectors for efficient similarity search. Similarly, Contriever [13] leverages contrastive learning to enhance semantic understanding. Advanced methods like E5 [48], BGE [29], and GTE [24] further utilize pre-trained or fine-tuned transformers for robust and efficient semantic retrieval.
@@
Long Term Memory Management. With the development of LLMs, the user-assistant conversation becomes longer and contains various topics [28, 5, 37, 32, 55, 56, 51, 5, 18], which introduces challenges for preserving the user’s long-term memory. Management in long-term memory often involves segmentation [33], summary [17], compression [3], forgetting and updating [1, 49, 59]. For example, several approaches [49, 42, 45, 26, 28] focus on generating memory summary as records to enable more accurate retrieval. Besides, some methods [33, 20, 52, 2] leverage compression techniques to reduce memory size while preserving essential information. The forgetting mechanisms [59, 14] address the need to remove obsolete or irrelevant memories while maintaining model performance. Some integrated methods [44, 31, 21] combine memory retention, update to enable personalized and contextually relevant long-term dialogue. However, Existing approaches typically focus on single-granularity segmentation strategies, such as session/turn, to organize and manage long-term memory. Whereas our work leverages multi-granular information for better adaptive memory selection.
@@
Structural Memory Management. Additionally, existing works employ structured paradigms for memory or knowledge base organization [4, 25, 55, 56, 51, 11, 57, 53]. HippoRAG [8, 9] builds entity-centric knowledge graphs inspired by hippocampal indexing theory [46], while Graph-CoT [15] and G-Retriever [12] integrate graph reasoning for interactive retrieval or generation, whereas LightRAG [7] and GraphRAG [6] optimize retrieval and summary via graph structures. MemTree and RAPTOR [39, 42] utilize recursive embedding, clustering, and summarization of text chunks to construct a hierarchical tree, while StructRAG [25] enhances reasoning by leveraging multiple structured formats. While existing methods focus on single-granularity memory modeling, lacking cross-granularity interactions. In contrast, our MemGAS, which leverages multi-granularity association and adaptive selection to construct consolidated memory structures and optimize retrieval efficiency.
@@ 5 Conclusion
Appendix A Datasets Statistics.
@@
Dataset LoCoMo-10 Long-MT-Bench+ LongMemEval-s LongMemEval-m
@@
  • LongMemEval-s [50] and LongMemEval-m [50] datasets are benchmarks designed to evaluate long-term memory in chat assistants. LongMemEval-s involves 50 sessions per question with an average of 115k tokens, making it a compact yet challenging dataset. LongMemEval-m, on the other hand, spans 500 sessions per question, resulting in avg. 1.5 million tokens pre conversations, offering a more extensive evaluation setting. Both datasets require AI systems to handle user-AI dialogues, dynamic memorization, and historical consistency.
  • LoCoMo [30] dataset evaluates long-term memory in AI through lengthy conversations, averaging 300 turns, 9,000 tokens, and up to 35 sessions. LoCoMo-10, a publicly available subset of LoCoMo, includes 10 high-quality, long conversations with 27.2 sessions and 20,000 tokens on average. The goal of LoCoMo is to evaluate long-context LLMs and RAG systems, which, while improving memory performance, still fall short of human capabilities—particularly in temporal reasoning—underscoring the challenges of understanding long-range dependencies.
  • Long-MT-Bench+ [33] dataset is reconstructed from MT-Bench+ [28] by incorporating long-range questions to address the limited QA pairs and short dialogues. It merges five consecutive sessions into a single long-term conversation, improving its suitability for evaluating memory mechanisms. The dataset features 11 conversations with an average of 4.9 sessions and 19194.8 tokens. Unlike LoCoMo-10, it focuses on user-AI interactions, excluding session dates and retrieval ground truth.
Table 6: QA performance.
Granularity GPT4o-J F1 BLEU Rouge1 Rouge2 RougeL BERTScore

### v2

@@ 3.2 Overall Results
We present the results for Question Answering and Retrieval in Table 1 and Table 2, respectively. We also compare performence of single-granularity and multi-granularity in Appendix C. Below, we provide analysis of these results.
Question Answering Results. As presented in Table 1, MemGAS consistently outperforms other methods across most datasets and evaluation metrics. Unlike Full History, which introduces noise by utilizing all historical context, MemGAS excels by effectively consolidating and retrieving only the most relevant memories. Other baselines, such as RecurSum and SeCom, although operating at specific granularities, lack the capability to integrate multi-granular associations between memory, resulting in suboptimal outcomes. Besides, methods like HippoRAG 2 and RAPTOR, while establishing connections between memory units, fail to construct multi-granular relationships and selection mechanisms effectively, limiting their performance. MemGAS performs better through its dynamic construction and adaptive router of multi-granular memory units and redundancy filtering, emphasizing the critical role of association and selection in memory management. Notably, these gains come with competitive efficiency: we keep average tokens and latency close to lightweight retrievers and below HippoRAG2 and A-Mem, achieving substantial improvements without incurring significant token or time overhead.
@@
MPNet (Song et al., 2020) 66.17 75.47 76.38 78.29 85.11 80.63
Contriever (Izacard et al., 2021) 71.06 79.72 81.28 82.47 90.00 84.29
RecurSum (Wang et al., 2025) 67.23 78.33 79.79 81.76 87.66 83.28
MPC (Lee et al., 2023) 60.00 70.90 68.09 73.27 80.00 76.59
SeCom (Pan et al., 2025) 71.06 80.88 80.43 83.08 89.15 85.11
HippoRAG 2(Gutiérrez et al., 2025) 75.53 85.44 84.68 87.32 91.28 88.73
@@
MPNet (Song et al., 2020) 37.02 48.68 45.74 52.51 61.28 56.55
Contriever (Izacard et al., 2021) 44.26 56.11 53.40 59.44 65.96 62.74
RecurSum (Wang et al., 2025) 23.19 32.10 31.70 36.99 43.62 41.40
MPC (Lee et al., 2023) 35.96 47.51 42.55 50.36 54.26 53.88
SeCom (Pan et al., 2025) 44.26 56.61 55.32 60.76 66.60 63.78
@@
LoCoMo
MPNet (Song et al., 2020) 45.92 47.71 53.98 51.79 68.58 56.88
Contriever (Izacard et al., 2021) 49.90 52.15 58.26 56.29 71.80 60.92
RecurSum (Wang et al., 2025) 47.23 48.99 59.01 54.58 74.97 60.07
MPC (Lee et al., 2023) 49.50 51.47 57.45 55.53 71.85 60.47
SeCom (Pan et al., 2025) 55.24 57.90 64.80 62.36 78.30 66.97
HippoRAG 2(Gutiérrez et al., 2025) 56.60 58.37 65.06 62.50 78.05 66.79
@@ 3.3 Ablation Study
The ablation study in Table 3 demonstrates the significance of each component in enhancing both QA and retrieval performance. Individually removing GMM, PPR, MA, or the Router results in consistent performance degradation, validating their essential contributions. Notably, the combined absence of all components leads to the most significant drop, with the F1 score plummeting from 20.38 to 13.78, and Recall@3 decreasing from 78.51 to 71.06. This highlights the importance of all modules in improving overall performance. Moreover, the latency introduced by these modules is minimal, with the highest latency increase being only 0.0191 seconds for QA and 0.0079 seconds for retrieval. This demonstrates that the proposed architecture achieves a remarkable balance between enhanced performance and computational efficiency.
Table 3: Ablation Study on LongMemeval-s for Gaussian Mixture Model (GMM), Personalized PageRank (PPR), Granularity Router and Memory Association (MA). The w/o MA setting is equivalent to w/o GMM and PPR. QA performance is evaluated using top 3 retrieved results. R@n means Recall@n. The ($\Delta$) represents the latency introduced by each module.
@@
GPT4o-J F1 RogueL Avg. Tokens Total Latency (s) R@3 R@5 R@10 Retrieval Latency (s)
MemGAS 60.20 20.38 19.47 8,829 2.5534 78.51 88.94 94.47 0.0239
w/o GMM 57.20 19.49 18.68 8,820 2.5506($\Delta$0.0028) 76.38 85.53 91.28 0.0232($\Delta$0.0007)
w/o PPR 56.60 19.76 18.85 8,772 2.5449($\Delta$0.0085) 75.96 85.96 90.64 0.0194($\Delta$0.0045))
w/o MA 56.80 17.69 19.00 8,734 2.5418($\Delta$0.0116) 74.89 85.74 91.49 0.0182($\Delta$0.0057)
w/o Router 56.60 18.88 18.62 8,763 2.5471($\Delta$0.0063) 75.53 85.74 92.34 0.0216($\Delta$0.0023)
w/o All 55.40 13.78 12.89 8,701 2.5343($\Delta$0.0191) 71.06 81.28 90.00 0.0160($\Delta$0.0079)
3.4 Detailed Comparison Analysis
@@
Comparison on different query types. The results in Figure 5 highlight performance across different query types. Our MemGAS consistently demonstrates superior performance, particularly excelling in multi-hop retrieval on LoCoMo and multi-session on LongMemEval-s. This suggests that memory association mechanism in MemGAS effectively identifies highly relevant sessions, enabling enhanced multi-hop reasoning and better integration of knowledge. Additionally, our method performs well across single-session, temporal reasoning, and knowledge update, demonstrating its strength in addressing both simple and complex query types. We also provide a case study for different query types in Appendix J.1, and comparison with other datasets and metrics in Appendix E.3.
Comparison on different Top-K retrieval. The results shown in Figure 4 demonstrate the performance of various Top-K retrieval settings. On the LoCoMo and LongMemEval-s datasets, our model consistently surpasses baseline methods in F1 scores as Top-K increases from 1 to 10. This trend highlights that retrieving more relevant context significantly enhances accuracy. Note that on LongMemEval-s dataset, while F1 scores initially improve at lower Top-K values, performance declines at higher Top-K levels, suggesting that the longer context introduces noise, which negatively impacts the model’s effectiveness.
@@ 4 Related Work
Retrieval Methods. Recent advancements in retrieval methods have greatly enhanced information retrieval performance(Robertson et al., 2009; Song et al., 2020; Santhanam et al., 2021). BM25 (Robertson et al., 2009), a classic sparse retrieval method, uses a probabilistic model to improve query relevance. In contrast, dense retrieval methods excel at capturing semantic similarity. For instance, DPR (Karpukhin et al., 2020) employs dual-encoder models to encode queries and documents into dense vectors for efficient similarity search. Similarly, Contriever (Izacard et al., 2021) leverages contrastive learning to enhance semantic understanding. Advanced methods like E5 (Wang et al., 2022), BGE (Luo et al., 2024), and GTE (Li et al., 2023) further utilize pre-trained or fine-tuned transformers for robust and efficient semantic retrieval.
@@
Long Term Memory Management. With the development of LLMs, the user-assistant conversation becomes longer and contains various topics (Chhikara et al., 2025; Li et al., 2025b; Wang & Chen, 2025; Lu et al., 2023; Du et al., 2024; Qian et al., 2025; Packer et al., 2023; Zhang et al., 2025; 2024a; Wu et al., 2025b; Du et al., 2024; Kirmayr et al., 2025), which introduces challenges for preserving the user’s long-term memory. Management in long-term memory often involves segmentation (Pan et al., 2025), summary (Kim et al., 2024), compression (Chen et al., 2025b), forgetting and updating (Bae et al., 2022; Wang et al., 2025; Zhong et al., 2024). For example, several approaches (Wang et al., 2025; Sarthi et al., 2024; Team, 2023; Liu et al., 2023a; Lu et al., 2023) focus on generating memory summaries as records to enable more accurate retrieval. Besides, some methods (Pan et al., 2025; Lee et al., 2024; Xu et al., 2023; Chen et al., 2025a) leverage compression techniques to reduce memory size while preserving essential information. The forgetting mechanisms (Zhong et al., 2024; Jia et al., 2024) address the need to remove obsolete or irrelevant memories while maintaining model performance. Some integrated methods (Tan et al., 2025; Ong et al., 2025; Li et al., 2024a) combine memory retention, update to enable personalized and contextually relevant long-term dialogue. However, existing approaches typically focus on single-granularity segmentation strategies, such as session/turn, to organize and manage long-term memory. Whereas our work leverages multi-granular information for better adaptive memory selection.
@@
Structural Memory Management. Additionally, existing works employ structured paradigms for memory or knowledge base organization (Chen et al., 2024; Li et al., 2024c; Zhang et al., 2025; 2024a; Wu et al., 2025b; He et al., 2024a; Zhang et al., 2024b; Xu et al., 2022; Rasmussen et al., 2025). HippoRAG (Gutiérrez et al., ; Gutiérrez et al., 2025) builds entity-centric knowledge graphs inspired by hippocampal indexing theory (Teyler & DiScenna, 1986), while Graph-CoT (Jin et al., 2024) and G-Retriever (He et al., 2024b) integrate graph reasoning for interactive retrieval or generation, whereas LightRAG (Guo et al., 2024) and GraphRAG (Edge et al., 2024) optimize retrieval and summary via graph structures. MemTree and RAPTOR (Rezazadeh et al., 2024; Sarthi et al., 2024) utilize recursive embedding, clustering, and summarization of text chunks to construct a hierarchical tree, while StructRAG (Li et al., 2024c) enhances reasoning by leveraging multiple structured formats. While existing methods focus on single-granularity memory modeling, they lack cross-granularity interactions. In contrast, our MemGAS, which leverages multi-granularity association and adaptive selection to construct consolidated memory structures and optimize retrieval efficiency.
@@ 5 Conclusion
Ethics Statement
This work adheres to ethical research practices by ensuring that all experiments and datasets used are publicly available and utilized in accordance with their licenses. The proposed MemGAS framework is designed to enhance conversational agents responsibly, with a focus on improving user experience while minimizing the risk of harm, such as generating misinformation. No sensitive or proprietary data was used during the research process, and the methodology prioritizes transparency and accountability in memory retrieval and usage.
Reproducibility Statement
We are committed to ensuring the reproducibility of our work. To this end, we have provided the code in an anonymous repository. The repository includes a well-documented README file with instructions to replicate our experiments. Additionally, we have detailed implementation specifics in this manuscript, and hyperparameter tuning is comprehensively described in Appendix F. Furthermore, we have included the complete set of prompts required for the experiments, ensuring that all components of our methodology can be accurately reproduced. We encourage the community to leverage these resources to build upon our work.
Appendix A Datasets Statistics
@@
Dataset LoCoMo Long-MT-Bench+ LongMemEval-s LongMemEval-m
@@
  • LongMemEval-s (Wu et al., 2025a) and LongMemEval-m (Wu et al., 2025a) datasets are benchmarks designed to evaluate long-term memory in chat assistants. LongMemEval-s involves 50 sessions per question with an average of 115k tokens, making it a compact yet challenging dataset. LongMemEval-m, on the other hand, spans 500 sessions per question, resulting in avg. 1.5 million tokens pre conversations, offering a more extensive evaluation setting. Both datasets require AI systems to handle user-AI dialogues, dynamic memorization, and historical consistency.
  • LoCoMo (Maharana et al., 2024) dataset evaluates long-term memory in AI through lengthy conversations, averaging 300 turns, 9,000 tokens, and up to 35 sessions. LoCoMo, the publicly available subset of the original paper, includes 10 high-quality, long conversations with 27.2 sessions and 20,000 tokens on average. The goal of LoCoMo is to evaluate long-context LLMs and RAG systems, which, while improving memory performance, still fall short of human capabilities—particularly in temporal reasoning—underscoring the challenges of understanding long-range dependencies.
  • Long-MT-Bench+ (Pan et al., 2025) dataset is reconstructed from MT-Bench+ (Lu et al., 2023) by incorporating long-range questions to address the limited QA pairs and short dialogues. It merges five consecutive sessions into a single long-term conversation, improving its suitability for evaluating memory mechanisms. The dataset features 11 conversations with an average of 4.9 sessions and 19194.8 tokens. Unlike LoCoMo, it focuses on user-AI interactions, excluding session dates and retrieval ground truth.
Appendix B Comparison of Methods Structure with Baselines
We compare the model structures of each baseline (provided in parentheses) and offer concise summaries to highlight the differences in their memory construction and retrieval mechanisms.
  • MPNet (Vector Base): Uses permuted language modeling with auxiliary positional information to model token dependencies while seeing full-sentence position signals.
  • Contriever (Vector Base): Trains dense retrievers via unsupervised contrastive learning for generalizable text embeddings, including cross-lingual retrieval.
  • MPC (Memory Pool): Composes LLM modules with few-shot prompting, chain-of-thought, and external memory to maintain long-term conversational consistency without fine-tuning.
  • RecurSum (Recursive Summary): Recursively generates and updates dialogue summaries as memory, using prior memory and new context to maintain coherence.
  • SeCom (Semantic Segment): Segments conversations into topically coherent units and applies compression-based denoising to build and retrieve from segment-level memory.
  • RAPTOR (Hierarchical Tree): Recursively embeds, clusters, and summarizes text to form a hierarchical summary tree, enabling retrieval at multiple abstraction levels.
  • HippoRAG2 (Knowledge Graph): Augments RAG with knowledge-graph traversal, integrating passages and online LLM reasoning for associative retrieval.
  • A-Mem (Memory Note): Implements a Zettelkasten-inspired memory: creates structured notes with attributes, indexes and links them, and updates existing notes as new ones arrive.
  • MemGAS (Ours; Multi-Granularity Memory): Builds multi-granularity memory units, associates them via Gaussian Mixture Models, and uses an entropy-based router plus LLM filtering for adaptive retrieval.
Table 5: QA performance of Single-Granularity and Multi-Granularity.
Granularity GPT4o-J F1 BLEU4 Rouge1 Rouge2 RougeL BERTScore